# MAV: A Multimodal, Multigenre, and Multipurpose Audio-Visual Academic Lecture Dataset

### v1

@@ Abstract
Publishing open-source academic video recordings is an emergent and prevalent approach to sharing knowledge online. Such videos carry rich multimodal information including speech, the facial and body movements of the speakers, as well as the texts and pictures in the slides and possibly even the papers. Although multiple academic video datasets have been constructed and released, few of them support both multimodal content recognition and understanding tasks, which is partially due to the lack of high-quality human annotations. In this paper, we propose a novel multimodal, multigenre, and multipurpose audio-visual academic lecture dataset (MAV), which has almost 367 hours of videos from five sources covering computer science, mathematics, and medical and biology topics. With high-quality human annotations of the spoken and written words, in particular high-valued name entities, the dataset can be used for multiple audio-visual recognition and understanding tasks. Evaluations performed on contextual speech recognition, speech synthesis, and slide and script generation tasks demonstrate that the diversity of MAV makes it a challenging dataset.
MAV: A Multimodal, Multigenre, and Multipurpose Audio-Visual Academic Lecture Dataset
Zhe Chen, Heyang Liu, Wenyi Yu, Guangzhi Sun, Hongcheng Liu Ji Wu, Chao Zhang🖂, Yu Wang🖂, Yanfeng Wang Shanghai JiaoTong University Tsinghua University University of Cambridge Department of Engineering Shanghai AI Laboratory
@@ 1 Introduction
The rapid progress of technology has brought numerous academic presentations and talks available on the web with open access Lev et al. (2019); Atri et al. (2021); Lee et al. (2023). These resources are particularly helpful to researchers as they contain rich specialised knowledge in auditory and visual modalities. With the development of neural-based AI systems, it is desired that AI could comprehend and process these multimodal information resources to assist scientists in accelerating their research. Abilities, including transcribing and generating speech for a presentation with the aid of slides, as well as presentation slide or script generation, are particularly useful for researchers to conduct investigations and prepare presentations.
As a natural data source containing multimodal content and knowledge, academic lectures have been explored to construct a series of datasets. Some datasets focus on the evaluation of the model’s ability to recognize multimodal content Dutta et al. (2018); Wang et al. (2023), others focus on promoting models to understand multimodal knowledge in academic videos Lev et al. (2019); Li et al. (2019); Atri et al. (2021); Lee et al. (2023). However, none of them has yet simultaneously paid attention to the model’s ability to recognize multimodal content and understand rich academic knowledge, which is partly due to their respective lack of adequate manual labelling. The homogeneity of holding two characteristics is critical to the implementation of end-to-end academic recognition and understanding systems.
To fill the gap, we release the Multimodal, Multigenre, and Multipurpose Audio-Visual Academic Lecture Dataset (MAV) as shown in Figure 1. MAV serves as a benchmark to evaluate the ability to perform multimodal perception and academic knowledge comprehension. The dataset contains selected academic lectures and presentation videos from multiple fields, including computer science, biomedical science, and mathematics. Each video is annotated with high-qualified speech transcriptions and OCR labels for both printed and handwritten characters, including complex mathematical formulae. In addition, academic papers for some videos are also provided to supplement knowledge more extensively.
Together with the dataset, three benchmark tasks are also proposed that reflect the perception and understanding of the multimodal information in those videos. This includes automatic speech recognition (ASR) with a particular focus on contextual ASR (CASR), spontaneous text-to-speech (TTS) synthesis, and slide and script generation (SSG). Representative baseline models are used to benchmark the performance for each task 111The dataset and the codes of benchmarks will be made available at [URL]. Through our experiments, it is observed that the existing models showcase limited performance in perceiving and understanding the multimodal content in this dataset, and the rich academic knowledge is not utilized effectively.
@@ 2 Related Work
LectureVideoDB Dutta et al. (2018) exploits frames from lecture videos to test the model’s ability to perform handwritten and scene text recognition. SlideSpeech Wang et al. (2023) enriches the audio-visual corpus with synchronized slides to provide additional textual information. It is worth noting that MAV dataset provides both OCR data that is fully manually labelled and speech data that combines multiple ASR system outputs and manual labelling, which is completely qualified to cover tasks in the aforementioned datasets.
@@
LectureBank Li et al. (2019) manually annotates the prerequisite relationships between course concepts to carry out prerequisite chain learning. GoogleI/O Chen et al. (2014) and LaRochelle Nguyen et al. (2014) both study video-level retrieval using presentation or lecture videos. AVIATE Atri et al. (2021) adopts multimodal information from academic presentation videos to conduct abstract summarization. Similarly, TalkSumm Lev et al. (2019) uses a scalable annotation method to construct large-scale paper summaries from corresponding video transcription. LPM Lee et al. (2023) introduces crossmodal retrieval and generation tasks around its aligned slides and spoken language. Our MAV dataset contains annotated video data for both slides and speeches, which implies little knowledge loss. The addition of paper documentation information also greatly complements the knowledge information. Furthermore, a challenging Slide and Script Generation task is developed to perform the inter-transformation of two kinds of academic materials.
@@ 3.1 Metadata
In OCR annotations, the coordinates and manually corrected text (formulae are in LaTeX format) of each block are provided in a single JSON file named ocr_result.json. Results of merging (detailed in Appendix A.1) are also provided.
@@ 3.2 Characteristics of Speech Data
As shown in Figure 2(a), MAV contains academic presentations and talk videos from three different fields including computer science, biomedical science, and mathematics. The Conference on Human Factors in Computing Systems (CHI) 2021 Presentations and ACM International Conference on Ubiquitous Computing (Ubi) 2020 Presentations focus on research in computer science. The National Institutes of Health Director’s Wednesday Afternoon Lectures (NIH) presents weekly science talks given by top biomedical scientists. The Introduction to the Principles and Practice of Clinical Research (IPP) teaches how to conduct clinical research effectively and safely. Oxford Mathematics (MLS) contains mathematics lectures. The total number of videos is 1113, with 366.9 hours of speech.
@@
MAV is labelled by combining multiple high-performing ASR systems together with manual labels. As human often fails to label rare terminologies, the assistance of a high-performing ASR is indispensable. The multimodal labelling pipeline is described in Section 4.2.
@@ 3.3 Characteristics of Visual Data
Table 4 shows the enrichment of our complex texts around which users can study to extract sophisticated semantic information. Furthermore, it can be noticed that the complex texts are primarily in the mathematics academic videos, which makes sense.
@@ 4.2.1 Step 1: Candidate Generation
The Microsoft transcriptions in written and spoken form are denoted as and respectively, and the Whisper transcriptions result in written form are denoted as (only the written form exists in the Whisper result).
@@ 4.2.2 Step 2: Audio Segmentation
To establish a uniform segmentation, re-segmentation is conducted using the word-level timestamps in . The split is allowed to be in silence and punctuation. And the duration of each segment is kept to 10 seconds or less. Detailed rules can be found in Appendix A.1.
@@ 4.2.4 Step 4: Candidate Combination
In Case I where is "approximately equal"666it forgives misalignments due to pauses and common words, as they are mainly short-pronounced, and and have high confidence in them. to , and will be directly adopted.
@@ 4.3.1 Step 1: Extensive Pre-Annotation
We conduct frame sampling for the video at 1 frame per second and use PaddleOCR777https://github.com/PaddlePaddle/PaddleOCR for large-scale text detection and recognition. PaddleOCR is selected for its completeness and high accuracy of the recognition results with little sticking between words.
@@ 4.3.3 Step 3: Manual Correction
For frames with simple text, corrections are directly executed by experienced researchers for the location and content of the OCR results. Note that blocks of presentation-irrelevant text such as URLs and email addresses are omitted. For frames with complex handwriting or formulae, we pre-label them using the Mathpix OCR API888https://mathpix.com/ocr. Then, the annotator performs careful corrections for the OCR result due to their complexity. The formulae are labelled in the format of in-line LaTeX. The details of manual labelling are shown in Appendix A.4. To control the quality of labelling, we sampled the results of 9 annotators individually and checked whether 120 of the annotated images had a high number of missing or incorrect items, with no more than 10 wrong items.
@@ 4.3.4 Step 4: Rule-Based Merging
To provide comprehensive semantic information, we devise three rules to perform merging:
@@ 5.1.2 Benchmark Systems
We adopt two widely-used benchmark systems for ASR and CASR, namely Attention-based Encoder-Decoder (AED) Chan et al. (2016) and RNN-Transducer (RNN-T) Graves (2012). Both systems employ a Conformer Gulati et al. (2020) encoder and the AED system is jointly trained and decoded with Connectionist Temporal Classification (CTC) Graves et al. (2006); Watanabe et al. (2017). For CASR, the Tree-Constrained Pointer Generator (TCPGen) Sun et al. (2023b) with GNN tree encodings Sun et al. (2023a) is used which employs a combination of symbolic prefix-tree search and a neural pointer generator for contextual biasing, detailed in Appendix B.1 and Appendix B.2.
| Model | Biasing List | Dev | Test | ||||
| CER () | WER () | BWER () | CER () | WER () | BWER () | ||
| CTC-Attention | - | 4.6 | 10.0 | 35.5 | 4.5 | 10.0 | 35.6 |
| RNN-T | - | 4.7 | 10.3 | 36.8 | 4.7 | 10.6 | 38.0 |
| Whisper-large-v3 | - | 3.9 | 7.1 | 19.4 | 3.9 | 7.2 | 18.9 |
| RNN-T + TCPGen | Random Distractors | 4.7 | 9.7 | 26.7 | 4.7 | 9.7 | 27.1 |
| RNN-T + TCPGen | OCR | 4.5 | 9.4 | 22.9 | 4.6 | 9.6 | 25.0 |
@@ 5.1.3 Experimental Results
In addition to the character error rate (CER) and word error rate (WER) for ASR, we use biasing WER (BWER) for CASR. BWER is defined as the total rare word errors divided by the total number of rare words, including insertions of rare words. Table 5 shows the experimental results. The AED system achieves 10.0% WER on both the Dev and Test sets. However, end-to-end models suffer from rare word recognition as reflected by the BWER where a more than three times increase in the error rate is observed comparing BWER to the overall WER for both AED and RNN-T systems. The results of Whisper-large-v3 Radford et al. (2023) are for reference only, as the Whisper-series model is involved in dataset construction and it only outputs the written form. Its massive training data and large size also result in incomparability with other systems.
For the CASR task, when the TCPGen is used without biasing, the experimental performance is roughly the same as RNN-T. When 1000 distractors are randomly selected from all rare words and are added to the biasing list, significant improvements in rare word recognition can be observed. When using OCR information together with rare words from the same topic to form the biasing list of each utterance during decoding, the recognition performance slightly deteriorates. This is because strongly related words are included in the biasing list, which may cause more confusion since words with related meanings often have similar pronunciations and word forms. When only OCR words are used, we achieve the best performance of the CASR task. In conclusion, by using TCPGen utilizing the OCR information, we achieve a relative BWER decrease of 37.8% and 34.2% on dev and test sets respectively.
@@ 5.2.1 Task Description
Spontaneous Text-to-Speech (TTS) Guo et al. (2021); Yan et al. (2021) focuses on generating speech with a natural and conversational session. Most TTS systems use clean speech. This form makes it easier to extract pronunciation features, while it places higher requirements on the quality of the training audio and often leads to a discernible gap between the synthesized output and authentic speech. We are devoted to using real speech data in MAV for the TTS system to produce speech that aligns more closely with natural conversational patterns, as it is spontaneous and has regulated pronunciations.
@@ 5.2.2 Benchmark Systems
We adopt a novel method as our spontaneous TTS baseline. MQTTS (multi-codebook vector quantized TTS) Chen et al. (2023) employs a two-stage training architecture. In the first stage, a quantizer is trained to discretize the raw waveform into a sequence of codes, minimizing quantization and reconstruction losses. A discriminator guides reconstruction via adversarial learning. In the second stage, the fixed quantizer is combined with a Transformer that performs autoregressive generation on the code sequence, conditioned on a speaker embedding for multi-speaker capability. More details can be found in Appendix B.3.

### v2

# M3AV: A Multimodal, Multigenre, and Multipurpose Audio-Visual Academic Lecture Dataset

@@ Abstract
Publishing open-source academic video recordings is an emergent and prevalent approach to sharing knowledge online. Such videos carry rich multimodal information including speech, the facial and body movements of the speakers, as well as the texts and pictures in the slides and possibly even the papers. Although multiple academic video datasets have been constructed and released, few of them support both multimodal content recognition and understanding tasks, which is partially due to the lack of high-quality human annotations. In this paper, we propose a novel multimodal, multigenre, and multipurpose audio-visual academic lecture dataset (M3AV), which has almost 367 hours of videos from five sources covering computer science, mathematics, and medical and biology topics. With high-quality human annotations of the slide text and spoken words, in particular high-valued name entities, the dataset can be used for multiple audio-visual recognition and understanding tasks. Evaluations performed on contextual speech recognition, speech synthesis, and slide and script generation tasks demonstrate that the diversity of M3AV makes it a challenging dataset111Project website: https://jack-zc8.github.io/M3AV-dataset-page.
M3AV: A Multimodal, Multigenre, and Multipurpose Audio-Visual Academic Lecture Dataset
Zhe Chen1, Heyang Liu1, Wenyi Yu2, Guangzhi Sun3, Hongcheng Liu1 Ji Wu2, Chao Zhang2🖂, Yu Wang1,4🖂, Yanfeng Wang1,4 1Department of Electronic Engineering, Shanghai JiaoTong University 2Department of Electronic Engineering, Tsinghua University 3University of Cambridge Department of Engineering 4Shanghai AI Laboratory {chenzhe2018,liuheyang,hongcheng_liu,yuwangsjtu,wangyanfeng622}@sjtu.edu.cn ywy22@mails.tsinghua.edu.cn, gs534@cam.ac.uk, {wuji_ee,cz277}@tsinghua.edu.cn
@@ 1 Introduction
The rapid progress of technology has brought numerous academic presentations and talks available on the web with open access Lev et al. (2019); Atri et al. (2021); Lee et al. (2022, 2023). These resources are particularly helpful to researchers as they contain rich specialised knowledge in auditory and visual modalities. With the development of neural-based AI systems, it is desired that AI could comprehend and process these multimodal information resources to assist scientists in accelerating their research. Abilities, including transcribing and generating speech for a presentation with the aid of slides, as well as presentation slide or script generation, are particularly useful for researchers to conduct investigations and prepare presentations.
As a natural data source containing multimodal content and knowledge, academic lectures have been explored to construct a series of datasets. Some datasets focus on the evaluation of the model’s ability to recognize multimodal content Dutta et al. (2018); Wang et al. (2023), others focus on promoting models to understand multimodal knowledge in academic videos Lev et al. (2019); Li et al. (2019); Atri et al. (2021); Lee et al. (2022, 2023). However, none of them has yet simultaneously paid attention to the model’s ability to recognize multimodal content and understand rich academic knowledge, which is partly due to their respective lack of adequate manual labelling. The homogeneity of holding two characteristics is critical to the implementation of end-to-end academic recognition and understanding systems.
To fill the gap, we release the Multimodal, Multigenre, and Multipurpose Audio-Visual Academic Lecture Dataset (M3AV) as shown in Figure 1. The dataset contains selected academic lectures and presentation videos from multiple fields, including computer science, biomedical science, and mathematics. Each video is endowed with highly qualified speech transcriptions and OCR labels for both printed and handwritten characters, including complex mathematical formulae. In addition, academic papers for some videos are also provided to supplement knowledge more extensively.
Together with the dataset, three benchmark tasks are also proposed that reflect the perception and understanding of the multimodal information in those videos. This includes automatic speech recognition (ASR) with a particular focus on contextual ASR (CASR), spontaneous text-to-speech (TTS) synthesis, and slide and script generation (SSG). Representative baseline models are used to benchmark the performance for each task. Through our experiments, it is observed that the existing models showcase limited performance in perceiving and understanding the multimodal content in this dataset, and the rich academic knowledge is not utilized effectively.
@@ 2 Related Work
LectureVideoDB Dutta et al. (2018) exploits frames from lecture videos to test the model’s ability to perform handwritten and scene text recognition. SlideSpeech Wang et al. (2023) enriches the audio-visual corpus with synchronized slides to provide additional textual information. It is worth noting that M3AV dataset provides both OCR data that is fully manually labelled and speech data that combines multiple ASR system outputs and manual labelling, which is completely qualified to cover tasks in the aforementioned datasets.
@@
LectureBank Li et al. (2019) manually annotates the prerequisite relationships between course concepts to carry out prerequisite chain learning. GoogleI/O Chen et al. (2014) and LaRochelle Nguyen et al. (2014) both study video-level retrieval using presentation or lecture videos. AVIATE Atri et al. (2021) adopts multimodal information from academic presentation videos to conduct abstract summarization. Similarly, TalkSumm Lev et al. (2019) uses a scalable annotation method to construct large-scale paper summaries from corresponding video transcription. LPM Lee et al. (2022, 2023) introduces crossmodal retrieval and generation tasks around its aligned slides and spoken language. Our M3AV dataset contains annotated video data for both slides and speeches, which implies little knowledge loss. The addition of paper documentation information also greatly complements the knowledge information. Furthermore, a challenging Slide and Script Generation task is developed to perform the inter-transformation of two kinds of academic materials.
@@ 3.1 Metadata
In OCR annotations, the coordinates and manually corrected text (formulae are in LaTeX format) of each block are provided in JSON files. Results of merging (detailed in Appendix A.1) are also provided.
@@ 3.2 Characteristics of Speech Data
As shown in Figure 2(a), M3AV contains academic presentations and talk videos from three different fields including computer science, biomedical science, and mathematics. The Conference on Human Factors in Computing Systems (CHI) 2021 Presentations and ACM International Conference on Ubiquitous Computing (Ubi) 2020 Presentations focus on research in computer science. The National Institutes of Health Director’s Wednesday Afternoon Lectures (NIH) presents weekly science talks given by top biomedical scientists. The Introduction to the Principles and Practice of Clinical Research (IPP) teaches how to conduct clinical research effectively and safely. Oxford Mathematics (MLS) contains mathematics lectures. The total number of videos is 1113, with 366.9 hours of speech.
@@
M3AV is labelled by combining multiple high-performing ASR systems together with manual labels. As human often fails to label rare terminologies, the assistance of a high-performing ASR is indispensable. The multimodal labelling pipeline is described in Section 4.2.
@@ 3.3 Characteristics of Visual Data
Significant numbers of pages and blocks in Table 4 show the enrichment of our complex texts. Furthermore, handwriting and formula texts which both contain lots of presentation content and academic knowledge have already been processed into standard text or LaTeX format and thus can be directly utilised by large language models.
@@ 4.2.1 Step 1: Candidate Generation
The Microsoft transcriptions in written and spoken form666“Spoken form” refers to the representation of language as it is actually spoken or pronounced, while “written form” refers to the representation of language in its written or textual format. are denoted as and respectively, and the Whisper transcriptions result in written form are denoted as (only the written form exists in the Whisper result).
@@ 4.2.2 Step 2: Audio Segmentation
Most ASR models expect that the input audio comes in the form of a relatively short duration, usually up to a few seconds in length. Therefore, we conduct re-segmentation using the word-level timestamps in . The split is allowed to be in silence and punctuation. And the duration of each segment is kept to 10 seconds or less. Detailed rules can be found in Appendix A.1.
@@ 4.2.4 Step 4: Candidate Combination
In Case I where is "approximately equal"777It forgives misalignments due to pauses and common words, as they are mainly short-pronounced, and and have high confidence in them. to , and will be directly adopted.
@@ 4.3.1 Step 1: Extensive Pre-Annotation
We conduct frame sampling for the video at 1 frame per second and use PaddleOCR888https://github.com/PaddlePaddle/PaddleOCR for large-scale text detection and recognition. PaddleOCR is selected for its completeness and high accuracy of the recognition results with little sticking between words.
@@ 4.3.3 Step 3: Manual Correction
For frames with simple text, corrections are directly executed by experienced researchers for the location and content of the OCR results. Note that blocks of presentation-irrelevant text such as URLs and email addresses are omitted. For frames with complex handwriting or formulae, we pre-label them using the Mathpix OCR API999https://mathpix.com/ocr. Then, the annotator performs careful corrections for the OCR result due to their complexity. The formulae are labelled in the format of in-line LaTeX. The details of manual labelling are shown in Appendix A.4. To control the quality of labelling, we sampled the results of 9 annotators individually and checked whether 120 of the annotated images had a high number of missing or incorrect items, with no more than 10 wrong items.
@@ 4.3.4 Step 4: Rule-Based Merging
To provide comprehensive semantic information, we devise three rules to perform merging101010As demonstrated in the top part of Figure 1, the recognition results of the general OCR model are in rows. It is required to merge these lines into paragraphs to provide complete semantic information.:
@@ 5 Benchmarks and Experiments
| Model | Biasing List | Dev | Test | ||||
| CER () | WER () | BWER () | CER () | WER () | BWER () | ||
| ASR | |||||||
| AED | - | 4.6 | 10.0 | 35.5 | 4.5 | 10.0 | 35.6 |
| RNN-T | - | 4.7 | 10.3 | 36.8 | 4.7 | 10.6 | 38.0 |
| Whisper-v3 | - | 3.8 | 7.1 | 19.4 | 3.8 | 7.2 | 18.9 |
| Wav2Vec2+CTC | - | 10.6 | 28.3 | 60.1 | 10.6 | 28.5 | 62.4 |
| Conformer+Trans. | - | 4.6 | 10.7 | 37.8 | 4.6 | 10.5 | 38.2 |
| Conformer+CTC | - | 4.4 | 11.1 | 40.6 | 4.3 | 11.1 | 42.2 |
| SeamlessM4T-v2 | - | 12.4 | 17.1 | 36.7 | 12.4 | 18.6 | 40.7 |
| CASR | |||||||
| RNN-T + TCPGen | Random Distractors | 4.7 | 9.7 | 26.7 | 4.7 | 9.7 | 27.1 |
| RNN-T + TCPGen | OCR | 4.5 | 9.4 | 22.9 | 4.6 | 9.6 | 25.0 |
@@ 5.1.2 Benchmark Systems
We adopt two widely-used benchmark systems for ASR and CASR, namely Attention-based Encoder-Decoder (AED) Chan et al. (2016) and RNN-Transducer (RNN-T) Graves (2012). Both systems employ a Conformer Gulati et al. (2020) encoder. The AED system is jointly trained and decoded with Connectionist Temporal Classification (CTC) Graves et al. (2006); Watanabe et al. (2017).
For ASR, we also introduce a variety of state-of-the-art pretrained ASR models. Note that the results of Whisper-v3 are for reference only, as the Whisper-series model is involved in dataset construction. The performance of the other open-source ASR models can be fairly evaluated, given that our videos are properly transcribed for the first time, avoiding the issue of training data contamination. These pretrained models are selected from the SpeechColab ASR Benchmark (EN)111111https://github.com/SpeechColab/Leaderboard and Hugging Face Community121212https://huggingface.co/models?pipeline_tag=automatic-speech-recognition&sort=trending which include Whisper-v3 Radford et al. (2023), Wav2Vec2+CTC Grosman (2021), Conformer+Transducer131313https://catalog.ngc.nvidia.com/orgs/nvidia/teams/nemo/models/stt_en_conformer_transducer_xlarge, Conformer+CTC141414https://catalog.ngc.nvidia.com/orgs/nvidia/teams/nemo/models/stt_en_conformer_ctc_xlarge and SeamlessM4T-v2 Barrault et al. (2023)151515https://huggingface.co/facebook/seamless-m4t-v2-large.
For CASR, the Tree-Constrained Pointer Generator (TCPGen) Sun et al. (2023b) with GNN tree encodings Sun et al. (2023a) is used which employs a combination of symbolic prefix-tree search and a neural pointer generator for contextual biasing. More details are illustrated in Appendix B.1 and Appendix B.2.
@@ 5.1.3 Experimental Results
In addition to the character error rate (CER) and word error rate (WER) for ASR, we use biasing WER (BWER) for CASR. BWER is defined as the total rare word errors divided by the total number of rare words, including insertions of rare words. Table 5 shows the experimental results. The AED and RNN-T systems achieve around 10.0% WER on both the Dev and Test sets. However, such end-to-end models suffer from rare word recognition as reflected by the BWER where a more than two times increase in the error rate is observed comparing BWER to the WER.
When 1000 distractors161616We select a certain number of significant words, which are the words in the bias list we aim to recognize accurately. When decoding, a particular utterance only includes a few of these words, while the rest will appear as distractors. The use of distractors follows Le et al.; Sun et al.. are randomly selected from all rare words and are added to the biasing list, significant improvements in rare word recognition can be observed. When using OCR information together with rare words from the same topic to form the biasing list of each utterance during decoding, the recognition performance slightly deteriorates. This is because strongly related words are included in the biasing list, which may cause more confusion since words with related meanings often have similar pronunciations and word forms. When only OCR words are used, we achieve the best performance of the CASR task. In conclusion, by using TCPGen utilizing the OCR information, we achieve a relative BWER decrease of 37.8% and 34.2% on dev and test sets respectively.
@@ 5.2.1 Task Description
Spontaneous Text-to-Speech (TTS) Guo et al. (2021); Yan et al. (2021) focuses on generating speech with a natural and conversational session. Most TTS systems use clean speech. This form makes it easier to extract pronunciation features, while it places higher requirements on the quality of the training audio and often leads to a discernible gap between the synthesized output and authentic speech. We are devoted to using real speech data in M3AV for the TTS system to produce speech that aligns more closely with natural conversational patterns, as it is spontaneous and has regulated pronunciations.
@@ 5.2.2 Benchmark Systems
Multi-codebook vector quantized TTS (MQTTS) Chen et al. (2023) is introduced as our spontaneous TTS baseline. It employs a two-stage training architecture. In the first stage, a quantizer is trained to discretize the raw waveform into a sequence of codes, minimizing quantization and reconstruction losses. A discriminator guides reconstruction via adversarial learning. In the second stage, the fixed quantizer is combined with a Transformer that performs autoregressive generation on the code sequence, conditioned on a speaker embedding for multi-speaker capability. More details can be found in Appendix B.3.